### v2

@@ 2.3 Time-dependent extension of the Ramo-Shockley theorem
The propagation of the signal through the resistive electrodes can be rigorously described by using an extended form of the Ramo-Shockley theorem for conductive media. The contribution of the material resistivity to the signal formation is included in the time evolution of the weighting potential, which is the solution of the Maxwell equations in the quasi-static limit [28].
@@
The current induced by an elementary charge , that is expressed by Eq. 2.3 for perfectly conductive electrodes, is now obtained through a convolution of the transport and propagation effects
@@
Calculating the time-dependent weighting potential via the Maxwell equations in the quasi-static limit constitutes a well-defined problem. Although computationally intensive because of the third-order PDE and of the complicated geometry, numerical solutions of Cauchy problems benefit from extensive research and development across various sectors, offering numerous methodologies. In the subsequent sections, we present a comparative analysis of results derived from COMSOL MultiPhysics® [30], employing Finite Element Methods (FEMs), and a spectral method that we have developed for this case.
2.3.1 Time-dependent weighting potential computation with Finite Element Method
The COMSOL MultiPhysics® software package implements a solver based on the Finite Element Method which allows to solve the Maxwell equations in the quasi-static limit using the Backward Differentiation Formula (BDF) method [39, 40]. The toolkit allows to solve field equations by modeling geometries made from nearly arbitrary shapes.
Finite Element Method (FEM) computations rely on the process of dividing the physical surfaces and volumes into a set of discrete two- or three-dimensional elements, known as meshing. These elements are designed to locally approximate the potential using second-order polynomials. A significant advantage of the meshing approach is that finer meshes can be designed near the electrodes, where the variations of the potential are expected to be more pronounced, while coarser meshes can be adopted in regions of slowly varying potential, thus optimizing the computational resources.
@@ 2.3.1 Time-dependent weighting potential computation with Finite Element Method
impose ideal grounding to all electrodes except the one under study, to which a voltage ramp is applied (to avoid the discontinuity of the Heaviside step function ), with a ramping time chosen much shorter than the reaction time of the resistive material;
@@
2.3.2 Time-dependent weighting potential computation with spectral method
@@
Additional boundary conditions are needed to constraint the voltage of the chosen readout electrode to . This is achieved by defining a finite number of points, named voltage pins, corresponding only to the electrode ends and defining as a weighted sum of the potential induced by point-like charges satisfying the boundary conditions on . This mechanism mimics the effect of an ideal voltage supplier injecting charge in a limited number of pre-defined point of the sensor as required to guarantee its set point.
@@
where and indicate a Fourier Transform and its inverse, respectively, and are implemented using the FFT module of the JAX library [41].
The weighting potential obtained with this method for the geometry of our test sensor using a mesh of 120x120x64 voxels is shown in Figure 7 at six different times.
@@ 2.4 Comparison of signal simulation methods
In the new simulation approach, the effect of resistance is modeled through time-dependent weighting potential maps, computed with either COMSOL MultiPhysics® or the spectral method. These maps are processed by Garfield++ implementing the convolution between the time evolution of the induced current due to the carrier transport and that due to the signal propagation as described in Equation 2.6.
2.4.1 Comparison of signals generated into the detector
@@
The current induced on the considered readout electrode using the maps obtained with the FEM implementation in COMSOL MultiPhysics® is compared to the simplified method representing the graphitic electrodes as discrete impedance networks in Figure 9. The comparison is performed using the same simulated incident particle, with identical trajectory and ionization energy deposits, and the same electrostatic field map determining the trajectories of the charge carriers. The two models differ significantly in the early stages of the signal development where the simple impedance network fails to capture the current induction close to the readout electronics input during the evolution of the time-dependent weighting field. Indeed, once the charge carriers have been all collected and the current is dominated by the delayed component due to the signal propagation through the electrodes, the predictions of the two models are consistent.
@@
The signals obtained from another incident particle using the FEM and the spectral methods are compared in Figure 9. The FEM signal, derived from a variable size mesh, is smoother than that obtained with the spectral method on a uniform voxel matrix. The spikes produced by the spectral method are indeed artifacts due to the rapid variation of the potential in the proximity of the electrodes which is not well represented by the fixed size voxels. This problem could be mitigated with a finer voxel pitch at the price of a considerably longer computing time. Nevertheless, the two signals agree very well and the computing artifacts are washed out once the raw signal is convoluted with any realistic readout electronics response function as discussed in the next section.
2.4.2 Effects of readout electronics and electronic noise
@@
In order to compare the simulation results with the data collected during the beam test described in Ref. [21], it is essential to properly take into account the effects due to the signal processing in the various steps of the readout. This is obtained by convolving the signal obtained from the detector simulation described in the previous sections with the transfer function of the readout electronics, and properly taking into account the corresponding electronic noise. For a full comparison, features of the data acquisition chain such as digitalization sampling frequency will also be considered.
@@
The readout front-end board employed in the beam test was developed at the University of Kansas (referred to as the "KU board" in the following) [42]. The KU board is based on a low cost two-stage common-emitter amplifier built with fast Si-Ge transistors. The board, originally developed for Ultra-Fast Silicon detectors, has been adapted to diamond sensors by increasing the gain of the amplifier to deal with the lower amount of charge released in diamond. This goes at the expenses of the timing performance which was not optimal for the 3D sensors used in the test.
@@
The output signals of the KU board were acquired through a 6 GHz digital oscilloscope at a sampling rate of 20 GS/s saving the waveforms on disk. Therefore, to properly simulate the full readout chain, the simulated detector signals were convolved with the KU board transfer function and interpolated at 20 GS/s. We considered the effect of the analog 6 GHz oscilloscope bandwidth negligible compared with the KU board bandwidth.
@@
Since the KU board has a non-trivial noise spectrum, in the simulation we empirically summed noise waveforms acquired with the oscilloscope in auto-trigger mode to the simulated waveforms. An example of waveform obtained through the whole simulation workflow is shown in Figure 10, where the results obtained with impedance network and time-dependent weighting potentials are superposed.
@@
At the end of the simulation flow discussed so far, both the experimental and simulated waveforms can be compared by extracting parameters like their amplitude, rise-time and delay with respect to a reference timing signal to measure the time resolution. The latter can be measured using for example a digital constant fraction discriminator algorithm, as described in detail in Ref. [43].
@@
The distributions of the simulated signal amplitudes and time markers, utilizing the time-dependent weighting potential as computed with the Finite Element Method (FEM) and spectral method, and obtained with the discrete impedance network are presented in Figure 11.
@@
Within the limits of the spectral method that must be refined as discussed in the previous sections, we consider the results obtained with the Finite Element and with the spectral methods consistent in the description of the signal amplitude spectrum. The simple discrete impedance description, besides being ad hoc, tends to slightly overestimate the signal amplitudes but is qualitatively consistent with the two previous methods. The time distribution of the signal is largely insensitive to the underlying induced current distributions, as it is dominated by the shaping effects of the readout electronics, and is only weakly dependent on the method employed to calculate the weighting potential. The results discussed above show that the weighting potentials obtained with the finite elements and the spectral methods give consistent and realistic results and we thus consider them sufficiently robust to allow a direct comparison with experimental data, as discussed in the next section.
@@ 3 Validation of simulation on beam-test data
To validate the simulation chain described above, we perform a comparison with data obtained during a beam test at CERN SPS in 2021 [21]. The tested device was a prototype 3D pixel matrix with m2 pitch and a geometry identical to that described in Section 2. The diameter of the electrodes, measured with an optical microscope from the shadow projected by the nano-structured graphitized material, is 12 m. As the electrodes terminate with only one end at the diamond crystal surface (see Figure 2), direct measurement of their resistance was not feasible. An indirect evaluation was therefore performed using specially fabricated graphitized through-columns, in which both ends were accessible. These reference electrodes were produced on the same diamond crystal, within the same fabrication run, and under identical process parameters as those adopted for the sensor. On this basis, the resistance of the sensor electrodes was estimated to be approximately 30 k. The electrode fabrication parameters were carefully optimized to lower as much as possible the resistance but the actual optical and lattice conditions may worsen the quality of the graphitization, making the measured 30 k more a lower limit than an expected value.
For the beam test, to enhance the active area, several pixels were connected in parallel into a strip of six cells and a comb-like arrangement of thirty cells by means of surface graphitic traces, as illustrated in Figure 12. To select the beam particles traversing the active area, we used a single-pixel silicon sensor with an active area of m2 [44] combined with a Micro-Channel Plate Photomultiplier Tube (MCP-PMT) which provided a trigger signal with a timing precision better than 10 ps [45]. The position of the silicon pixel relative to the diamond was adjusted using a piezoelectric linear stage in order to allow a scan of the diamond matrix active area. The signals were read out and acquired through the electronics chain described in Section 2.4.2.
@@
Figure 14 shows the amplitude distributions obtained from the simulation and the test beam data. The two distributions exhibit a similar structure and share common features. In the simulation, a peak near 0 V is observed, which can be attributed to three main categories of events: events where both end points of the particle trajectory lie within the readout electrode, events with the final end-point outside the simulation volume, and events occurring in proximity to either of these categories. Beyond this peak, the simulated distribution displays the expected Landau shape. Between the two peaks of the distribution, there is a number of events associated with the presence of inclined tracks, where part of the charge is created outside the cell volume. A Landau fit of the main peak yields a most probable amplitude for the simulated amplitudes of
@@
The experimental amplitude distribution presents three similar components: an initial peak near 0 V, primarily due to low-amplitude signals from particles traversing the graphite columns, charge sharing with adjacent pixels, and electronic noise; a valley, again associated with inclined tracks; a Landau peak, whose most probable value is found at . The 10% discrepancy with respect to the prediction from simulation is consistent with an increased detector capacity due to the multiple electrodes connected to form the strip, but could also be due to an underestimation of the electrode resistance in the simulation setup, or even to a sub-fluctuation of the amplifier gain due to fabrication tolerances on the transistors.
@@
The time resolution of both the simulated and experimental device is obtained from the distribution of the signal delays with respect to a reference timing signal. For the simulated events the reference time is conventionally set to zero while for the real data this is provided by the MCP-PMT detector. The observed time distribution is not Gaussian and shows a significant tail at large relative times due to several effects which however can be of different origin. In data, the long relative time tail is populated by low amplitude signals coming for example from the noise distribution tails or due to charge sharing between adjacent pixels. In simulation instead, a significant number of events in the distribution upper tail are due to artifacts coming from the inaccurate modeling of charge carrier motion inside the electrodes.
@@
This The relative weight of the tail can thus be reduced by rejecting low amplitude events. In Figure 14 simulation and data are compared for signals with an amplitude larger than 60 mV. The two distributions are still slightly asymmetric although the central Gaussian core has a significantly larger weight. To properly take into account the residual distribution asymmetry, a fit is performed with a Crystal Ball function which features a Gaussian core and an exponential tail. The Gaussian core can be conventionally used to estimate the time resolution which is measured to be ps in the simulated events and ps in the experimental data. The discrepancy is largely due to the differences in the signal amplitude spectrum discussed above.
@@
Considering the differences between the simulated system and the experimental device as well as the large uncertainty on the resistance of the graphitized electrodes, the agreement between data and simulation can be considered satisfactory. For example repeating the simulation choosing a resistance for the electrodes of 60 k, the predicted amplitude peak becomes mV and the time resolution ps. Hence, we considered the simulation procedure validated and ready to serve as a valuable tool to guide the optimization of detector performance through targeted modifications in sensor fabrication techniques.

### v3

@@ 2.4 Time-dependent extension of the Ramo-Shockley theorem
The propagation of the signal through the resistive electrodes can be rigorously described by using an extended form of the Ramo-Shockley theorem for conductive media. The contribution of the material resistivity to the signal formation is included in the time evolution of the weighting potential, which is the solution of the Maxwell equations in the quasi-static limit [Rigler:2e004jh].
@@
The current induced by an elementary charge , that is expressed by Eq. 2.3 for perfectly conductive electrodes, is now obtained through a convolution of the transport (process 2) and propagation effects (process 3)
@@
Calculating the time-dependent weighting potential via the Maxwell equations in the quasi-static limit constitutes a well-defined problem. Although computationally intensive because of the third-order PDE and of the complicated geometry, numerical solutions of Cauchy problems benefit from extensive research and development across various sectors, offering numerous methodologies. In the subsequent sections, we present a comparative analysis of results derived from COMSOL MultiPhysics® [Janssens:2890572], employing FEM solvers (), and a spectral method () that we have developed for this case.
2.4.1 Time-dependent weighting potential computation with a FEM solver ()
The COMSOL MultiPhysics® software package implements a solver based on FEMs which allows to solve the Maxwell equations in the quasi-static limit using the Backward Differentiation Formula (BDF) method [Curtiss:1952, Gear:1967]. The toolkit allows to solve field equations by modeling geometries made from nearly arbitrary shapes.
Computations with FEMs rely on the process of dividing the physical surfaces and volumes into a set of discrete two- or three-dimensional elements, known as meshing. These elements are designed to locally approximate the potential using second-order polynomials. A significant advantage of the meshing approach is that finer meshes can be designed near the electrodes, where the variations of the potential are expected to be more pronounced, while coarser meshes can be adopted in regions of slowly varying potential, thus optimizing the computational resources.
@@ 2.4.1 Time-dependent weighting potential computation with a FEM solver ()
impose ideal grounding to all electrodes except the one under study, to which a voltage ramp is applied. We impose ideal grounding to all electrodes except the one under study, to which a voltage ramp is applied. Here, is a smoothed version of the Heaviside step function, with a sub-femtosecond rise time chosen much shorter than the response time of the resistive material;
@@
2.4.2 Time-dependent weighting potential computation with spectral method ()
@@
Additional boundary conditions are needed to constraint the voltage of the chosen readout electrode to . This is achieved by defining a finite number of points, named voltage pins, corresponding only to the electrode ends and defining as a weighted sum of the potential induced by point-like charges satisfying the boundary conditions on . This mechanism mimics the effect of an ideal voltage supplier injecting charge in a limited number of pre-defined point of the sensor to guarantee its setting conditions.
@@
where and indicate a Fourier Transform and its inverse, respectively, and are implemented using the FFT module of the JAX library [jax2018github].
The weighting potential obtained with this method for the geometry of the sensor under study using a mesh of voxels is shown in Figure 7 at six different times.
@@ 2.5 Comparison of signal simulation methods
In the new simulation approach, the effect of resistance is modeled through time-dependent weighting potential maps, computed with either COMSOL MultiPhysics® () or the spectral solver (). These maps are processed by Garfield++ implementing the convolution between the time evolution of the induced current due to the carrier transport and that due to the signal propagation as described in Equation 2.6.
@@
The current induced on the considered readout electrode using the maps obtained with the FEM implementation in COMSOL MultiPhysics® () is compared to the simplified method representing the graphitic electrodes as discrete impedance networks () in Figure 9. The comparison is performed using the same simulated 180 GeV/ pion, with identical trajectory and ionization energy deposits, and the same electrostatic field map determining the trajectories of the charge carriers. The two models differ significantly in the early stages of the signal development, where the simple impedance network fails to capture the current induction close to the readout electronics input during the evolution of the time-dependent weighting field. Indeed, once the charge carriers have been all collected and the current is dominated by the delayed component due to the signal propagation (process 3) through the electrodes, the predictions of the two models are consistent.
The disagreement between the predictions in the first part of the signal, corresponding to the time interval where process 2 and process 3 coexist, raises concerns about the reliability of at lower electrode resistances. Process 3 becomes negligible in the limit of metallic electrodes, where it is effectively instantaneous, and the agreement between models is ensured by the identical treatment of process 2. However, there is no evidence that the simplified model correctly captures the coupled dynamics of processes 2 and 3 when the resistance is finite. This limitation becomes particularly relevant when considering optimization strategies for detector geometry or readout electronics, as the early signal region, where the discrepancy is most pronounced, plays an important role in determining the final time resolution.
@@
The signals obtained from another incident particle using the FEM () and the spectral methods () are compared in Figure 9. The FEM signal, derived from a variable size mesh, adapted to the geometry, is smoother than that obtained with the spectral method on a uniform voxel matrix. The spikes produced by the spectral method are indeed artifacts due to the rapid variation of the potential in the proximity of the electrodes which is not well represented by the fixed size voxels. This problem could be mitigated with a finer voxel pitch at the price of a considerably longer computing time. Nevertheless, the two signals agree very well and the computing artifacts are washed out once the raw signal is convoluted with any realistic readout electronics response function as discussed in the next section.
2.6 Effects of readout electronics and electronic noise
@@
In order to compare the simulation results with the data collected during the beam test described in Ref. [Anderlini:2022exf], it is essential to properly take into account the effects due to the signal processing in the various steps of the readout. This is obtained by convolving the signal obtained from the sensor simulation described in the previous sections with the transfer function of the readout electronics, and properly taking into account the corresponding electronic noise. For a full comparison, features of the data acquisition chain such as digitalization sampling frequency will also be considered.
@@
The readout front-end board employed in the beam test was developed at the University of Kansas (referred to as the "KU board" in the following) [Minafra:2017nqc]. The KU board is based on a low cost two-stage common-emitter amplifier built with fast Si-Ge transistors. The board, originally developed for Ultra-Fast Silicon detectors, has been adapted to diamond sensors by increasing the gain of the amplifier to deal with the lower amount of charge released in diamond. This goes at the expenses of the timing performance which was not optimal for the 3D sensors used in the test.
@@
The output signals of the KU board were acquired through a 6 GHz digital oscilloscope at a sampling rate of 20 GS/s, saving the waveforms on disk. Therefore, to properly simulate the full readout chain, the simulated detector signals were convolved with the KU board transfer function and interpolated at 20 GS/s. We considered the effect of the analog 6 GHz oscilloscope bandwidth negligible compared with the KU board bandwidth.
@@
Since the KU board has a non-trivial noise spectrum, in the simulation we empirically summed noise waveforms acquired with the oscilloscope in auto-trigger mode, with the sensor connected and polarized, to the simulated waveforms. An example of waveform obtained through the whole simulation workflow is shown in Figure 10, where the results obtained with impedance network and time-dependent weighting potentials are superposed.
@@
At the end of the simulation flow discussed so far, both the experimental and simulated waveforms can be compared by extracting parameters like their amplitude, rise-time and delay with respect to a reference timing signal to measure the time resolution. The latter can be measured using for example a digital constant fraction discriminator algorithm, as described in detail in Ref. [chiara].
@@
The distributions of the simulated signal amplitudes and time markers, utilizing the time-dependent weighting potential as computed with the FEM solver () and the spectral method (), and obtained with the discrete impedance network () are presented in Figure 11.
@@
Within the limits of the spectral method that must be refined as discussed in the previous sections, we consider the results obtained with and as consistent in the description of the signal amplitude spectrum. The simple discrete impedance description, besides being ad hoc, tends to slightly overestimate the signal amplitudes but is qualitatively consistent with the two previous methods. The time distribution of the signal is largely insensitive to the underlying induced current distributions, as it is dominated by the shaping effects of the readout electronics. Consistent results are obtained regardless of whether the weighting potential is computed using spectral or FEM solvers.
We conclude that both and provide predictions consistent with the heuristic in the regime where it is expected to be reliable, and consistent with each other, especially once the amplification and shaping effects (process 4) are taken into account. Hence, we consider and as sufficiently robust to allow a direct comparison with experimental data, as discussed in the next section.
@@ 3 Validation of simulation on beam-test data
To validate the simulation chain described above, we perform a comparison with data obtained during a beam test at CERN SPS in 2021 [Anderlini:2022exf]. The tested device was a prototype 3D pixel matrix with m2 pitch and a geometry identical to that described in Section 2. The diameter of the electrodes, measured with an optical microscope from the shadow projected by the nano-structured graphitized material, is 12 m. As the electrodes terminate with only one end at the diamond crystal surface (see Figure 2), direct measurement of their resistance was not feasible. An indirect evaluation was therefore performed using specially fabricated graphitized through-columns, in which both ends were accessible. These reference electrodes were produced on the same diamond crystal, within the same fabrication run, and under identical process parameters as those adopted for the sensor. On this basis, the resistance of the sensor electrodes was estimated to be approximately 30 k. The electrode fabrication parameters were carefully optimized to lower as much as possible the resistance but the actual optical and lattice conditions may worsen the quality of the graphitization, making the measured 30 k more a lower limit than an expected value.
For the beam test, to enhance the active area, several pixels were connected in parallel into a strip of six cells and a comb-like arrangement of thirty cells by means of surface graphitic traces, as illustrated in Figure 12. To select the beam particles traversing the active area (or equivalently, to veto particles at the boundaries of the sensor or beyond), we used a single-pixel silicon sensor with an active area of m2 [Anderlini:2020ffm] combined with a Micro-Channel Plate Photomultiplier Tube (MCP-PMT) which provided a trigger signal with a timing precision better than 10 ps [Garau:2023fyo]. The position of the silicon pixel relative to the diamond was adjusted using a piezoelectric linear stage in order to allow a scan of the diamond matrix active area. The signals were read out and acquired through the electronics chain described in Section 2.6.
@@
Figure 14 shows the amplitude distributions obtained from the simulation and the test beam data. The two distributions exhibit a similar structure and share common features.
@@
In the simulation, a peak near 0 V is observed, which can be attributed to three main categories of events: events where both end points of the particle trajectory lie within the readout electrode, events with the final end-point outside the simulation volume, and events occurring in proximity to either of these categories. Beyond this peak, the simulated distribution displays the expected Landau shape. Between the two peaks of the distribution, there is a number of events associated with the presence of inclined tracks, where part of the charge is created outside the cell volume. A Landau fit of the main peak yields a most probable amplitude for the simulated amplitudes of
@@
The experimental amplitude distribution presents three similar components: an initial peak near 0 V, primarily due to low-amplitude signals from particles traversing the graphite columns, charge sharing with adjacent pixels, and electronic noise; a valley, again associated with inclined tracks; a Landau peak, whose most probable value is found at .
The 10% discrepancy with respect to the prediction from simulation is consistent with an increased detector capacity due to the multiple electrodes connected to form the strip, but could also be due to an underestimation of the electrode resistance in the simulation setup, or even to a sub-fluctuation of the amplifier gain due to fabrication tolerances on the transistors.
@@
The time resolution of both the simulated and experimental device is obtained from the distribution of the signal delays with respect to a reference time signal. For the simulated events, the reference time is conventionally set to zero, while for the real data this is provided by the MCP-PMT detector. The observed time distribution is not Gaussian and shows a significant tail at large relative times, due to several effects which, however, can be of different origin.
In data, the long relative time tail is populated by low amplitude signals coming for example from the noise distribution tails or due to charge sharing between adjacent pixels. In simulation, instead, a significant number of events in the upper tail of the distribution are due to artifacts arising from the inaccurate modeling of charge carrier motion inside the electrodes.
@@
The relative weight of the tail can thus be reduced by rejecting low-amplitude events. In Figure 14, simulation and data are compared for signals with an amplitude larger than 60 mV. The two distributions are still slightly asymmetric although the central Gaussian core has a significantly larger weight. To properly take into account the residual distribution asymmetry, a fit is performed with a Crystal Ball function which features a Gaussian core and an exponential tail. The Gaussian core can be conventionally used to estimate the time resolution which is measured to be ps in the simulated events and ps in the experimental data. The discrepancy is largely due to the differences in the signal amplitude spectrum discussed above.
@@
Considering the differences between the simulated system and the experimental device, as well as the large uncertainty on the resistance of the graphitized electrodes, the agreement between data and simulation can be considered satisfactory. For example, repeating the simulation with a resistance of 60 k for the electrodes, the predicted amplitude peak becomes mV. The corresponding time resolution is ps. Hence, we considered the simulation procedure validated and ready to guide the optimization of detector performance through targeted modifications in sensor fabrication techniques.